Find the linearizationL(x,y,z) of the function f(x,y,z) =sin(xy)/z at the point (pi/2,1,1).
1
Expert's answer
2020-10-06T18:33:01-0400

Given, "f:\mathbb{R}^3\rightarrow\mathbb{R}" to be a real valued map "(x,y,z)\mapsto\frac{\sin(xy)}{z}" , now we have to find Linearization of "f(x,y,z)" at "(a,b,c)=(\pi/2,1,1)"


Since, from Taylor expansion we get


"L(x,y,z)=f(a,b,c)+\sum_{x,yz}(x-a)f_x(a,b,c)"

Now,

"f_x(x,yz)=\frac{y}{z}\cos(xy)\implies f_x(\pi/2,1,1)=0\\
f_y(x,y,z)=\frac{x}{z}\cos(xy)\implies f_y(\pi/2,1,1)=0\\
f_z(x,y,z)=-\frac{1}{z^2}\sin(xy)\implies f_z(\pi/2,1,1)=-1"

Thus,

"L(x,y,z)=1+(x-\pi/2)\cdot 0+(y-1)\cdot 0+(z-1)\cdot(-1)\\
\implies L(x,y,z)=-z+2"